OpenStudy (anonymous):

Find the distance between the points with the given coordinates (-4, 9) and (1, -3).

OpenStudy (anonymous):

distance formula is \[\sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\]

OpenStudy (anonymous):

how do that

OpenStudy (ivettef365):

you can use website, enter the points and will calculate the distance

OpenStudy (anonymous):

\[\sqrt{(1-(-4)^{2}+(-3-9)^{2}}\] first subtract 1 from -4 and -3 from 9

OpenStudy (anonymous):

\[\sqrt{(5)^{2}+(-12)^{2}}\] now square 5 and -12

OpenStudy (anonymous):

\[\sqrt{25+144}=\sqrt{169}=13\]

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

no problem-0 :)

OpenStudy (anonymous):

message me

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

k

OpenStudy (ivettef365):

@ElShortie100 did you try doing it on the website ?

OpenStudy (anonymous):

@ivettef365 thats helps alot thanks alot

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

yea